OpenStudy (anonymous):

Did I do this right? Please check my work, somebody? On January 1, 2010, Chessville has a population of 50,000 people. Chessville then enter a period of population growth. Its population increases 7% each year. On the same day, Checkersville has a population of 70,000 people. Checkersville starts to experience a population decline. Its population decreases 4% each year. During what year will the population of Chessville first exceed that of Checkersville?

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

HELP ME

OpenStudy (anonymous):

PLS SHOW STEPS IM DYING

OpenStudy (anonymous):

AHHHHHH

OpenStudy (anonymous):

show me what you have so far

OpenStudy (anonymous):

OK HOLD UP

OpenStudy (anonymous):

Y=A*B^x 50,000(1+0.07)^t=50,000(1.07)^t 70,000(1-0.04)^t=70,000(0.96)^t

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

help????

OpenStudy (anonymous):

@Fransisco_w

OpenStudy (anonymous):

i thought u had something for a bit

OpenStudy (anonymous):

50000*7%= 3500 70000*4%= 2800 50000+3500y=70000-2800y --> figure it out from here

OpenStudy (anonymous):

y=year

OpenStudy (anonymous):

?

OpenStudy (anonymous):

AHHH HELP I SKIPPED &TH AND *TH GRADE IS TOO HARD

OpenStudy (anonymous):

7th

OpenStudy (anonymous):

8th

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

HELP HELP

OpenStudy (anonymous):

alright ivwill work through it

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

*STARES AT COMPUTER SCREEN IN SUSSPENSE*

OpenStudy (anonymous):

first subtract 50000 from 70000 and then 2800y to 3500y to get 6300y=20000 y=3.17 final answer

OpenStudy (anonymous):

O_O

OpenStudy (anonymous):

OOOHHHHH MAKES SO MUCH SENSE THANK YOU GREAT AND POWERFUL ONE

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

how did you get 2800 and 3500

OpenStudy (anonymous):

@Fransisco_w

OpenStudy (anonymous):

@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w@Fransisco_w

OpenStudy (anonymous):

@Fransisco_w @Fransisco_w @Fransisco_w @Fransisco_w @Fransisco_w @Fransisco_w @Fransisco_w

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

@dragons1415

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (anonymous):

@TheGh0st

OpenStudy (anonymous):

@bohotness

OpenStudy (bohotness):

wEEEEEEEEEE

OpenStudy (anonymous):

D:

OpenStudy (bohotness):

*falls in question &*dusk selfo ff * what do you need herlp with i got your back

OpenStudy (anonymous):

Did I do this right? Please check my work, somebody? On January 1, 2010, Chessville has a population of 50,000 people. Chessville then enter a period of population growth. Its population increases 7% each year. On the same day, Checkersville has a population of 70,000 people. Checkersville starts to experience a population decline. Its population decreases 4% each year. During what year will the population of Chessville first exceed that of Checkersville?

OpenStudy (anonymous):

Y=A*B^x 50,000(1+0.07)^t=50,000(1.07)^t 70,000(1-0.04)^t=70,000(0.96)^t

OpenStudy (bohotness):

*sits in the question yes its right XD

OpenStudy (anonymous):

first subtract 50000 from 70000 and then 2800y to 3500y to get 6300y=20000 y=3.17 final answer

OpenStudy (anonymous):

where did the 2800y and 3500y com from?

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (bohotness):

okay

OpenStudy (anonymous):

@xapproachesinfinity com on ur my last hope

OpenStudy (anonymous):

smartest guy i know

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (bohotness):

o.o *leaves questoin and telps to someeone else*

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):

@Directrix ok ur the second smartest guy i know

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

me need kitty

OpenStudy (bohotness):

hi MEOW KITTY HERE